OpenStudy (anonymous):

3x+7y=12 6x-ky=6 for which of the following values for k will the system of equations above has no solution ? -14 -7 0 7 14

OpenStudy (anonymous):

Substitute each possible value into the equations to see if they work

OpenStudy (anonymous):

After you substitute a value in, solve the system. example: 6x - (-14)y = 6 3x + 7y = 12 I'd multiply - 2(3x + 7y = 12 ) then add it to the top equation. Does that help